OpenStudy (anonymous):

A mathematics journal has accepted 14 articles for publication. However, due to budgetary restraints only 9 articles can be published this month. How many ways can the journal editor assemble 9 of the 14 articles for publication?

OpenStudy (anonymous):

\[\dbinom{14}{9}=\dbinom{14}{5}=\frac{14\times 13\times 12\times 11\times 10}{5\times 4\times 3\times 2}\]

OpenStudy (anonymous):

apparently it is 2002 http://www.wolframalpha.com/input/?i=14+choose+9